OpenStudy (anonymous):

The legs of an isosceles triangle are 5 units long, and the altitude to the base is 4 units long find the length of the base.

OpenStudy (anonymous):

Convert the isosceles triangle into 2 right triangles by dropping a line from vertex to base. You now have a right triangle with hypotenuse = 5 and a leg =4. \[C^2=A^2+B^2\] will let you find half of the base length. C = 5 A = 4 B = 1/2 length of base

OpenStudy (anonymous):



OpenStudy (anonymous):

the length of the base is 6 then

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

That's what I have